OpenStudy (anonymous):

the answer a fraction??? 6(2n + 3) - 4n + 32 = 90

OpenStudy (anonymous):

12n+18-4n+32=90 8n=90-18-32 = 40 n=5

OpenStudy (anonymous):

cyter is correct n = 5